Ticket Number: HEA-NOC/20020104-01 Ticket Status: UPDATE
Ticket Type: Unscheduled Ticket Source: HEA-NOC
Ticket Scope: Line Site/Line: HEAnet <==> NUI
Ticket Owner: [log in to unmask] Problem Fixer: HEA-NOC
Ticket Opened: Problem Start: 20020104 04:52:38 UTC
Ticket Closed: Problem End: 20020104 06:18:45 UTC
See times edited below: From router logs 04:52:38 OSPF Neighbour down, Line protocol down,line protocol up, 06:18:45 OSPF neighbor up.
Attempted to contact NUI but their office is closed until Mon 7th Jan.
Contacted Eircom to do alarm check on (DS 0831547)from 04:30 to 06:30, ref number is 314.
NUI Link to HEAnet
Time to Fix: